### v2

# Gyro-based Neural Single Image Deblurring

Abstract
In this paper, we present GyroDeblurNet, a novel single image deblurring method that utilizes a gyro sensor to effectively resolve the ill-posedness of image deblurring. The gyro sensor provides valuable information about camera motion during exposure time that can significantly improve deblurring quality. However, effectively exploiting real-world gyro data is challenging due to significant errors from various sources including sensor noise, the disparity between the positions of a camera module and a gyro sensor, the absence of translational motion information, and moving objects whose motions cannot be captured by a gyro sensor. To handle gyro error, GyroDeblurNet is equipped with two novel neural network blocks: a gyro refinement block and a gyro deblurring block. The gyro refinement block refines the error-ridden gyro data using the blur information from the input image. On the other hand, the gyro deblurring block removes blur from the input image using the refined gyro data and further compensates for gyro error by leveraging the blur information from the input image. For training a neural network with erroneous gyro data, we propose a training strategy based on the curriculum learning. We also introduce a novel gyro data embedding scheme to represent real-world intricate camera shakes. Finally, we present a synthetic dataset and a real dataset for the training and evaluation of gyro-based single image deblurring. Our experiments demonstrate that our approach achieves state-of-the-art deblurring quality by effectively utilizing erroneous gyro data.
1 Introduction
Blur caused by camera shakes severely degrades image quality as well as the performance of various computer vision tasks such as classification, object detection, and semantic segmentation. To overcome the image degradation caused by blur, single image deblurring has been extensively studied for decades [6, 12, 52, 32, 48, 17, 14]. Recently, various DNN-based approaches have been proposed and have shown significant performance improvements over classical approaches [31, 53, 44, 23, 24, 51, 5, 4]. Nonetheless, they still fail on images with large blur due to the severe ill-posedness of deblurring, i.e., multiple different blur trajectories could lead to similar-looking blurry images.
To address the ill-posedness, several attempts have been made to exploit gyro sensors with which most smartphones are nowadays equipped. As gyro data are correlated with motion of the camera mounted on a smartphone, classical approaches based on blur models use gyro data to guide blur kernel estimation [17, 40, 41, 11, 14, 29]. However, their performances are limited by their restrictive blur models that cannot effectively handle noise, nonlinear camera response functions, saturated pixels, moving objects, and so on. With the advance of deep learning, DNN-based approaches have also been proposed to exploit gyro sensors [30, 25, 16]. Thanks to the remarkable capability of DNNs and gyro sensors, they show better performance compared to classical gyro-based approaches and recent DNN-based approaches without gyro sensors.
However, recent DNN-based approaches that utilize gyro data are still limited in handling real-world blurred images. These approaches assume that the gyro data can accurately represent the blur in the blurred image, but this assumption does not hold in practice. First, gyro data from mobile devices such as smartphones usually have a significant amount of noise caused by various sources [19, 10, 17]. Second, gyro sensors measure the angular velocity around their center. Thus, to exploit gyro sensors, the recent gyro-based methods assume that the gyro sensor and the camera share the same center, and the camera is shaken only by rotational motions. However, in practice, the positions of the camera and the gyro sensor are different, and the camera shakes are caused by more complex motions.Third, real-world blurred images may have moving objects with different blur trajectories that cannot be captured by gyro sensors. All these factors make the motion information encoded in the gyro data inconsistent with the blur in the blurred image, and make it difficult to exploit the gyro data in real-world image deblurring. In this paper, we refer to such inconsistency between the gyro data and blur as gyro error.
In this paper, we propose GyroDeblurNet, a novel gyro-based single image deblurring approach that can produce high-quality deblurring results in the presence of gyro error. To address the gyro error, GyroDeblurNet adopts a carefully designed network architecture that contains a novel gyro refinement block and a novel gyro deblurring block. The gyro refinement block refines gyro data with error using the blur information in the input image. On the other hand, the gyro deblurring block removes blur in the input image using the refined gyro data. To further compensate for gyro error, the gyro deblurring block also exploits the blur information in the input image as well as the input gyro data. Despite the proposed blocks, training a neural network with erroneous gyro data is difficult as the network could be trained to simply ignore the input gyro data. To resolve this issue, we also propose a curriculum-learning-based training strategy [2].
To handle complex real-world camera shakes, we also present the camera motion field, a novel gyro data embedding scheme that can represent complex camera motions. Real-world blurry images often exhibit complex blur trajectories due to random camera shakes, which often occur when shooting a still shot with a long exposure time. Nonetheless, recent gyro-based approaches [30, 25, 16] assume smooth camera motions and represent camera motions using one or two motion vector per pixel or a couple of homographies for the entire image, which leads to low-quality deblurring results. To resolve this issue, our camera motion field is designed to represent complex camera motions using multiple vectors per pixel while maintaining a comparable memory footprint exploiting the spatially smooth nature of camera shakes.
For the training and evaluation of gyro-based deblurring, previous methods [30, 25, 16] rely on synthetic datasets. However, these datasets do not accurately reflect the noise and blur in real-world blurred images as shown by Rim et al. [35]. Additionally, their gyro data do not reflect real-world camera shakes that occur when capturing still-shot images either. Gyro data of previous methods were generated by random sampling [54] or obtained from the Visual-Inertial dataset [38]. However, the Visual-Inertial dataset is originally developed for visual odometry and SLAM, so its camera motions are collected from constantly moving cameras, which differ from camera shakes of still-shot images.
For training and evaluation of gyro-based deblurring for real-world blurred images, we propose two datasets: GyroBlur-Synth and GyroBlur-Real. GyroBlur-Synth is a synthetic dataset for both training and evaluation. The dataset consists of sharp and blurred image pairs with their corresponding gyro data. The gyro data are acquired from a smartphone device to reflect the real-world camera shakes. The blurred images are synthetically generated using the acquired gyro data and the blur synthesis process of RSBlur [34] for realistic blur. GyroBlur-Real is a real-world blur dataset for qualitative evaluation with no ground-truth sharp images. GyroBlur-Real provides 100 real-world blurred images paired with gyro data collected from a smartphone.
We validate the performance of GyroDeblurNet on both synthetic and real-world datasets. Our experiments demonstrate that GyroDeblurNet clearly outperforms state-of-the-art deblurring methods both quantitatively and qualitatively. Our contributions can be summarized as follows:
  • We propose a novel gyro-based single image deblurring approach, GyroDeblurNet. Extensive quantitative and qualitative evaluations show that GyroDeblurNet outperforms existing single image deblurring methods.
  • To handle gyro error, we introduce a carefully designed network equipped with a novel gyro refinement block and a novel gyro deblurring block. We also develop a novel gyro data embedding scheme to represent complex camera shakes and propose a curriculum learning-based training scheme.
  • We present synthetic and real image-gyro paired datasets with realistic camera motions, GyroBlur-Synth and GyroBlur-Real, for training and evaluating gyro-based single image deblurring methods.
2 Related Work
A variety of DNN-based single image deblurring approaches have recently been proposed [31, 53, 44, 23, 24, 51, 5, 4, 47, 45, 21]. To push the limit of single image deblurring, they introduce various network architectures and training schemes. Nevertheless, single image deblurring is still a challenging problem due to its high ill-posedness, and they still fail on images with large blur.
To resolve the ill-posedness of deblurring, there have been attempts to leverage inertial measurement sensors such as the gyro sensor and accelerometer as they provide valuable information on the rotational and translational motion of a camera, respectively. Joshi et al. [17] and Hu et al. [14] exploit the gyro sensor and accelerometer to estimate spatially-varying blur kernels. On the other hand, Park and Levoy [11] and Mustaniemi et al. [29] use only the gyro sensor to estimate blur kernels, because blur kernel estimation using accelerometer measurements requires estimating the depth of a scene, the device orientation to compensate the gravity effect, and the initial velocity of the device, which is often error-prone in practice. Moreover, rotational motion can be assumed as the dominant factor of camera shake blur as discussed by Whyte et al. [48].
In the deep learning era, there have been a few works that exploit gyro data. Mustaniemi et al. [30] propose a DNN-based approach that uses gyro data for the first time. They introduce a U-Net-based architecture [36] that takes a concatenation of a blurred image and the gyro data as input, which is also adopted by Lee et al. later [25]. Ji et al. [16] present a deformable-convolution-based network module for handling gyro data, which computes deformable convolution offsets from input gyro data to transform convolution filters according to the blur direction. However, these methods assume that the gyro data is accurate, neglecting potential errors. Moreover, they assume smooth camera motions and model camera motions using a motion vector field consisting of either one [30] or two motion vectors [16] per pixel, or a couple of homographies [25] representing the beginning and end position of the camera during the exposure time. Unfortunately, such restrictive assumptions do not hold for real-world blurred images with complex camera shakes as discussed in Sec. 1.
Utilizing gyro data by converting it into the form of motion vectors for single image deblurring can be considered as a non-blind deconvolution approach. While non-blind deconvolution methods have been studied for decades [33, 49, 22, 18, 26], they are not robust to blur kernel errors and produce unsatisfactory results with erroneous blur kernels. To overcome such limitation, several works adopted hand-crafted priors such as sparse prior [15], hyper-Laplacian prior [46, 22] and deep residual prior [43]. However, they may fail to utilize erroneous blur kernels where such prior assumptions do not hold.
3 GyroDeblurNet
The goal of GyroDeblurNet is to estimate a sharp deblurred image $S$ from an input blurred image $B$ and its corresponding gyro data $G$ possibly with large errors. Specifically, we define $G$ as a sequence of gyro data such that $G=\left\{g_{0},\cdots,g_{T-1}\right\}$, where $g_{t}$ is the $t$-th gyro data consisting of three angular velocities during the exposure time. $T$ is the number of gyro data, which is proportional to the exposure time. For example, $T=20$ if the exposure time is 0.1 seconds, and the sampling rate of the gyro sensor is 200 Hz. Instead of directly using $G$, we devise a novel gyro data embedding scheme named camera motion field to effectively handle complex camera shakes of an arbitrary exposure duration. In the rest of this section, we explain the gyro data embedding scheme based on the camera motion field, network architecture, and training strategy of GyroDeblurNet in detail.
Figure 1: Camera motion field construction. (a) Computing homography and warped pixel coordinates. (b) Constructing camera motion field by stacking motion vectors.
Figure 2: Network architecture of GyroDeblurNet.
Figure 3: Detailed architectures of the gyro refinement block and gyro deblurring block.
3.1 Gyro Data Embedding
GyroDeblurNet first converts input gyro data sequence $G$ into a camera motion field $\mathcal{V}$ before feeding it into the subsequent neural network to handle temporally complex camera shakes of an arbitrary length $T$. The conversion process is illustrated in Fig. 1. Our camera motion field has a fixed channel size $2M$ where $M$ is a hyperparameter, regardless of the length of the input gyro data to allow it to be fed to convolutional neural networks. To achieve this, we first resample the input gyro data sequence $G$ of an arbitrary length $T$ to $M+1$ samples using cubic-spline interpolation. We denote the resampled sequence as $G^{\prime}=\left\{g^{\prime}_{0},\cdots,g^{\prime}_{M}\right\}$ where $g^{\prime}_{m}$ is a resampled gyro sample. The hyperparameter $M$ is set to an even number so that we have the same number of gyro data before and after the temporal center of the exposure duration.
Then, assuming that the camera and the gyro shares the same center and there exists neither off-center rotational nor translational motions, we integrate the angular velocities and obtain $M+1$ camera orientations $\theta_{m}$, where $m=0,\cdots,M$, including the camera orientations at the beginning and end of the exposure. Note that, despite our restrictive assumption on the camera motion, GyroDeblurNet can successfully perform deblurring thanks to its robustness to gyro errors. For each $\theta_{m}$, we compute a homography $H_{m}$ as $H_{m}=KR(\theta_{m})K^{-1}$ where $K$ is the camera intrinsic matrix and $R(\theta_{m})$ is the rotation matrix corresponding to $\theta_{m}$. Based on the camera orientation that we chose as a reference, there can be multiple sets of homography $H_{m}$ each of which leads to shifted vector field. To prevent unnecessary shift after deblurring, we assume $\theta_{M/2}$ at the temporal center to be $(0,0,0)$ so that $H_{M/2}$ is an identity matrix when computing $\theta_{m}$.
Finally, from the obtained homographies, we compute a camera motion field $\mathcal{V}$, which illustrates how each pixel is blurred. Specifically, we define a camera motion field $\mathcal{V}$ as a tensor of size $W/{s}\times H/{s}\times 2M$ where $W$ and $H$ are the width and height of the input blurred image, and $s$ is a scaling factor. Then, at each spatial position $(x,y)$ in $\mathcal{V}$, we compute the warped coordinates of $(x,y)$ by $H_{m}$ for all $m$ and compute the difference between temporally consecutive coordinates to obtain $M$ motion guidance vectors at each position. By stacking the vectors for all the spatial positions, we construct $\mathcal{V}$. In our experiments, we set $M=8$ and $s=2$.
Our camera motion field is motivated by the blur field proposed by Mustaniemi et al. [30] and the motion guidance vector (MGV) field introduced by Ji et al. [16]. Unlike previous methods that rely on only one or two vectors to represent camera shakes at each pixel, our camera motion field offers a more sophisticated representation. By adjusting the hyperparameter $M$, it can effectively capture temporally intricate camera shake patterns. Furthermore, our approach leverages the spatial smoothness of camera shakes, utilizing the hyperparameter $s$. As a result, it can provide a detailed representation of complex motions while maintaining a comparable memory footprint.
3.2 Network Architecture
The network architecture of GyroDeblurNet is illustrated in Fig. 2. GyroDeblurNet takes a blurred image $B$ and a camera motion field $\mathcal{V}$ as input and estimates a deblurred image $S$. The network consists of two modules: an image deblurring module and a gyro module. The image deblurring module takes a blurred image $B$ and performs deblurring with the aid of gyro data. It then produces a residual $R$ that is added back to $B$ to produce $S$. The image deblurring module adopts a U-Net architecture [36] and the NAFBlock [4] as its basic building block. The image deblurring module adopts gyro deblurring blocks in its bottleneck to perform deblurring using gyro features from the gyro module.
On the other hand, the gyro module takes a camera motion field $\mathcal{V}$ and refines it with the aid of image features from the image deblurring module. The gyro module consists of a convolution layer to embed an input camera motion field $\mathcal{V}$ into the feature space, and gyro refinement blocks. The gyro refinement blocks refine gyro features with the aid of image features from the encoder of the image deblurring module. The gyro module progressively refines gyro features through gyro refinement blocks and strided convolution layers, and obtains a gyro feature of the same spatial size as the feature maps of the bottleneck in the image deblurring module. In the following, the gyro refinement block and the gyro deblurring block, which are the two key components of GyroDeblurNet, are explained in detail. For a detailed network architecture including feature dimensions, refer to the supplementary material.
Figure 4: The input camera motion field passes a convolutional layer to produce differently refined camera motion candidates in different channels. The candidates are then weighted by computing similarity with the blur information of the blurred image.
Gyro refinement block
The structure of the gyro refinement block is illustrated in Fig. 3. The gyro refinement block takes a gyro feature computed from the input camera motion field $\mathcal{V}$ as input. Then, it refines the gyro feature using the channel attention mechanism [13] with the aid of an image feature from the encoder of the image deblurring module. Specifically, the gyro refinement block takes an image feature from the image deblurring module as an additional input, and concatenates it with the input gyro feature. It then computes channel-wise attention weights through global average pooling and $1\times 1$ convolution layers, and refines the gyro feature using the obtained channel-wise attention weights. Finally, it applies a convolution layer to further refine the gyro feature.
The intuition behind the refinement process by the gyro refinement block is illustrated in Fig. 4. The camera motion information in the input camera motion field $\mathcal{V}$ may contain a significant amount of error. From such an erroneous camera motion information, the convolution layer at the beginning of the gyro module produces a gyro feature consisting of multiple camera motion candidates at different channels by introducing different perturbations to the input camera motion information. Then, the gyro refinement block adaptively selects the channels of the gyro feature that are consistent with the blur information encoded in the image feature using the channel attention weights computed from both gyro and image features [13].
Gyro deblurring block
The structure of the gyro deblurring block is shown in Fig. 3. The gyro deblurring block performs deblurring of the image feature with the gyro feature from the gyro module. The input image $B$ may have spatially-variant blur caused by camera shakes and moving objects. To effectively remove such spatially-varying blur in the image feature, the gyro deblurring block adopts the deformable convolution operation [7, 55], whose offsets are computed based on both image and gyro features. As the gyro feature provides accurately refined information on the camera shakes, and the image feature provides information on the blur in $B$ including object motion blur, the gyro deblurring block can successfully handle both camera shakes, and object motion blur.
To further remove the remaining spatially-variant blur in the deblurred image feature, the gyro deblurring block applies a spatial attention block to the deblurred image feature following Xu et al. [50]. The deblurred feature is then further refined through a NAFBlock and a convolution layer. While deformable convolution has also been employed by EggNet [16] to deblur images using gyro data, EggNet applies it without accounting for real-world gyro errors. This oversight leads to subpar results for real-world images, as will be shown in Sec. 5. In contrast, our approach incorporates network modules, a training strategy, and data specifically designed to accommodate real-world gyro errors.
3.3 Training Strategy
Despite the gyro refinement block and gyro deblurring block, it is still challenging to train a deblurring network with erroneous gyro data. Naïvely training the network with erroneous gyro data can lead the network to ignore the gyro data and simply rely only on the blur information in the blurred image. Thus, to train our network to fully utilize input gyro data, we introduce a training strategy based on curriculum learning [2], which has shown superior performance in various learning-based approaches that deal with erroneous data [37, 39, 42].
Our training strategy initially trains the network with error-free gyro data, and gradually increases error in the training data. To this end, we first assume that for each blurred image $B$, its corresponding error-free gyro data $G_{GT}$ is given as well as its ground-truth sharp image $S_{GT}$. From $G_{GT}$, we compute a noise-free camera motion field $\mathcal{V}_{clean}$ and a noisy camera motion field $\mathcal{V}_{noisy}$. We generate $\mathcal{V}_{noisy}$ by randomly perturbing the rotational center and adding noise to the gyro data $G_{GT}$. We then compute a blended camera motion field $\mathcal{V}_{\alpha}=(1-\alpha)\cdot\mathcal{V}_{clean}+\alpha\cdot\mathcal{V}_{noisy}$ where $\alpha$ is a blending parameter. Finally, we train our network with a PSNR loss [4, 3], which is defined as:
$\mathcal{L}_{PSNR}=10\cdot\log_{10}(\|f(B,\mathcal{V}_{\alpha})-S_{GT}\|^{2}+\epsilon)$ (1)
where $f(B,\mathcal{V}_{\alpha})$ represents a deblurred output of GyroDeblurNet and $\epsilon$ is a small constant to prevent the negative infinity, which is set to $81e-$ 8 in our experiments. We initially set $\alpha$ to 0, and gradually increase it to $1$ during training. For more details including the noisy camera motion field generation and the scheduling of $\alpha$, refer to the supplementary material.
4 GyroBlur Dataset
Figure 5: Examples of the GyroBlur-Synth and GyroBlur-Real datasets. In (a), the top and bottom rows show ground-truth images and the corresponding blurry images, respectively.
4.1 GyroBlur-Synth
GyroBlur-Synth provides 14,600 and 640 synthetically blurred images of size $720\times 1280$ for the training and test sets. The dataset also provides ground-truth sharp images, ground-truth clean gyro data. To faithfully mimic real-world blurred images, the dataset covers various sources of image degradation, such as camera shakes, moving objects, noise, and saturation. The dataset also includes gyro error data to produce realistic erroneous gyro data from the clean gyro data. Fig. 5 (a) shows examples of GyroBlur-Synth.
To build a dataset with realistic camera motions, we implemented an Android application that records gyro data on a mobile phone. We used a Samsung Galaxy S22 smartphone for collecting gyro data with the sampling rate of 200 Hz. We collected two sequences of gyro data for generating training and test datasets with realistic hand-shake motions. The gyro data sequences are recorded for 195 sec. and 60 sec., and have 39,116 and 12,065 samples, respectively. We then used the gyro data to synthesize blurry images. To this end, we used sharp frames of the 4KRD dataset [8]. To generate a blurry image, we first randomly sampled a sequence of $10$ consecutive gyro data corresponding to the exposure time of $1/20$ seconds. We then interpolated them by eight times to synthesize a continuous blur trajectory, and computed homographies from the interpolated gyro data. Finally, we warped the sharp image using the homographies, and blended them to obtain a blurry image.
To create synthetic blurred images with moving objects that have different blur trajectories, we used a simple method. We select a single instance from the COCO dataset [27] and assume it moves at a constant speed in a blurry image. We randomly choose its position and the direction and speed of motion. We then add the object to the warped sharp images before blending them during the blur synthesis process described earlier. This method may seem simple, but our experiment shows that it helps train our model to better handle spatially-variant motion blur from moving objects.
To reflect real-world noise and saturated pixels, we also adopt the blur synthesis pipeline of RSBlur [34]. Specifically, we synthesize blur in the linear sRGB space following the aforementioned process. Then, we synthesize saturated pixels, convert the image into the camera RAW space, synthesize noise, and convert the noisy RAW image to the linear sRGB image. We refer the readers to [34] for the detailed blur synthesis process. The blur synthesis pipeline of RSBlur requires camera parameters such as a color correction matrix and a noise distribution. To this end, we use the camera parameters of a Samsung Galaxy S22 ultra-wide camera, and the noise distribution estimated from it.
The gyro error data includes gyro noise and randomly perturbed rotation centers for generating noisy camera motion fields. To obtain realistic gyro noise, we measured the gyro noise distribution from a stationary smartphone (Samsung Galaxy S22) placed on a table. Gyro noise is modeled as a Gaussian distribution for each rotation axis independently. For more details on the dataset, we refer the readers to the supplementary material.
Our synthetic dataset generation approach requires minimal device-specific information, including a small amount of gyro data sequences and parameters. This approach minimizes the need for laborious manual data collection. Specifically, we can collect a sufficient amount of gyro data by holding a smartphone in hand only for a couple of minutes, and by placing a smartphone on a table only for a few seconds. The camera parameters required for the RSBlur pipeline can also be easily obtained by a simple calibration process with a couple of images captured by a target camera. Note also that our synthetic dataset approach process does not explicitly model the relative position of the gyro sensor with respect to the camera module, which is information that is hard to access. Instead, our approach implicitly models it as part of rotational center error. Despite this, our experiments show that our approach can effectively remove blur.
4.2 GyroBlur-Real
GyroBlur-Real provides 100 real-world blurry images and their corresponding gyro data for qualitative evaluation. The dataset was collected using a Samsung Galaxy S22. The blurry images are of size $4080\times 3060$, and provided in both raw DNG and JPEG formats. The dataset also provides metadata such as timestamps of the images and gyro data, camera intrinsic parameters, exposure times and ISO parameters. Fig. 5 (b) shows examples of GyroBlur-Real.
5 Experiments
For evaluation, we implemented our network using PyTorch. We trained our network on randomly cropped $256\times 256$ image patches from the GyroBlur-Synth dataset for 300 epochs with a batch size of 16. We used the Adam optimizer [20] with $\beta_{1}=0.9$ and $\beta_{2}=0.999$. The learning rate was initially set to 0.0001 and reduced to 1e-7 using the cosine annealing scheduler [28]. We measured the computation times of all the models on a GeForce RTX 3090 GPU.
One well-known issue in using gyro data is that the timestamps of the gyro sensor and the camera are not synchronized. Fortunately, several approaches that can successfully synchronize the gyro sensor and camera have been proposed [9]. Thus, in our work, we assume that the input gyro data are already synchronized for the exposure time of the input image. GyroBlur-Real dataset was collected using an open-source Android application [1] that supports synchronization between the gyro and camera based on gyro-based clock synchronization [9].
In the remainder of this section, we present the performance evaluation of GyroDeblurNet. We evaluate the performance of our method using the test set of GyroBlur-Synth and GyroBlur-Real. Additional experiments and examples including visualization of the effect of gyro refinement and an analysis on the robustness to gyro errors are included in the supplementary material.
Table 1: Quantitative comparison on the test of GyroBlur-Synth. The inference times of the models that use gyro data include both gyro data embedding (e.g., camera motion field construction) and deblurring processes.
  Gyro Models PSNR SSIM Param. (M) MACs (G) Time (s)
X MPRNet [51] 25.03 0.7081 20.13 10927 0.998
NAFNet [3] 25.06 0.7085 17.11 227 0.106
Uformer-B [47] 25.72 0.7334 50.88 2143 1.061
Stripformer [45] 25.93 0.7398 19.71 2397 1.367
FFTformer [21] 26.01 0.7481 16.56 1894 1.879
O DeepGyro [30] 23.78 0.6649 31.03 769 0.068
EggNet [16] 25.49 0.7266 6.34 1102 0.071
Ours 27.28 0.7803 16.31 262 0.130
Figure 6: Qualitative comparison on GyroBlur-Synth. The red curves overlaid on the blurred images visualize the input gyro data.
Figure 7: Qualitative comparison on GyroBlur-Real. The red curves overlaid on the blurred images visualize the input gyro data.
Comparison with state-of-the-art methods
We first compare GyroDeblurNet with state-of-the-art single image deblurring methods that use gyro data  [30, 16] and those that do not [51, 3, 47, 45, 21]. We evaluate the performance of all the models on the test set of GyroBlur-Synth with gyro errors. The results are presented in Tab. 1. All the models in the table were trained with GyroBlur-Synth using their own hyperparameter settings.
As shown in Tab. 1, GyroDeblurNet outperforms all the previous non-gyro-based methods by a large margin. Especially compared to Uformer-B [47], Stripformer [45] and FFTformer [21], which adopt transformers, our method requires a significantly smaller amount of computation, and a shorter computation time, while achieving more than 1 dB higher PSNR. Moreover, our method outperforms all the previous gyro-based methods by a large margin, while ours requires a comparable model size and a much smaller computational overhead.
Qualitative comparisons on GyroBlur-Synth and GyroBlur-Real are shown in Fig. 6 and in Fig. 7, respectively. As shown in the figures, the previous non-gyro-based methods fail to restore sharp details due to severe blur. Moreover, despite the input gyro data, the previous gyro-based methods also fail to produce sharp details as they are not robust to gyro errors. On the other hand, our method successfully restores sharp images as our method can effectively exploit gyro data with errors.
Figure 8: Camera motion field visualization on a GyroBlur-Synth image and a GyroBlur-Real image. (a) camera motion field visualization on a GyroBlur-Synth image. (b) Our result of (a). (c) A GyroBlur-Real image. (d) Gyro errors. (e) Our result of (c).
Gyro errors
As discussed in Sec. 1, real-world gyro data cannot accurately represent the camera motion due to gyro noise and missing information such as rotational centers. Fig. 8 shows examples of such gyro errors in our datasets where camera motion fields are visualized as red curves. As shown in the figure, the camera motion information in the camera motion fields in both GyroBlur-Synth and GyroBlur-Real do not perfectly match the blur trajectories in the blurred images due to gyro errors. Nonetheless, GyroDeblurNet successfully restores sharp images for both synthetic and real-world input images.
Table 2: Results on GyroBlur-Synth with two models trained with noisy and noise-free gyro data.
    Gyro data Noisy Noise-free
  PSNR 27.28 24.94
  SSIM 0.7803 0.7112
Effect of training with noisy gyro data
Our training data includes gyro errors to generalize our model to real-world gyro errors. Here we demonstrate the effect of the synthetically added gyro errors. Fig. 9 shows that the model trained with noisy gyro data successfully removes blur from a real-world blurred image with gyro errors, while the one trained with noise-free gyro data fails. Tab. 2 presents a quantitative comparison on GyroBlur-Synth. As shown in the table, training with noisy gyro data substantially improves the deblurring quality by more than 2 dB, proving that including gyro errors in the training data is crucial for the training of gyro-based deblurring.
Figure 9: Effect of training with noisy gyro data. (a) Real-world blurred image. (b) Model trained with noisy gyro data. (c) Model trained with noise-free gyro data.
Moving objects
Another source of gyro error is moving objects with different blur trajectories as discussed in Sec. 1. Fig. 10 shows a real-world example where the input blurry image has a moving object with different blur trajectory. Due to severe blur, the result of FFTformer [21], the state-of-the-art non-gyro-based method, still has remaining blur in both background and moving object. EggNet [16] also fails to deblur the image as it does not consider neither moving objects nor gyro errors. In contrast, our method successfully restores sharp details both on the moving object and the background thanks to its robustness to gyro errors.
Figure 10: Results of a real-world blurred image with a moving object.
Table 3: Effect of camera motion field hyperparameter $M$.
  $M$ 2 4 8 16
PSNR 25.71 26.80 27.28 27.32
SSIM 0.7706 0.7651 0.7803 0.7811
Camera motion field hyperparameter $M$
We investigate the impact of the hyperparameter $M$ that determines the temporal resolution of the camera motion field. To verify the effect of $M$, we compare the deblurring performance of different values of $M$. As shown in Tab. 3, the deblurring quality improves gradually as $M$ increases. This highlights the importance of gyro data representation that can capture real-world intricate camera motions, and also explains the limited performance of previous gyro-based approaches [16, 30]. Fig. 11 also demonstrates the effect of increasing the number of vectors to handle complex real-world motion blur. However, increasing $M$ to 16 does not yield a significant performance gain while it doubles the memory consumption, indicating that $M=8$ is sufficient to represent the camera shakes in most of the images in the test set of GyroBlur-Synth.
Figure 11: Effect of increasing $M$ for real-world complex motion blur. (a) Real-world blurry image with complex blur. (b) Result with $M=2$. (c) Result with $M=8$.
Table 4: Ablation study on the components of GyroDeblurNet.
  Model PSNR SSIM
(a) Deblurring with no gyro data 24.90 0.6997
(b) Gyro refine. w/o image features 26.64 0.7584
(c) Deform. conv. using only gyro features 26.68 0.7588
(d) GyroDeblurNet 26.94 0.7667
(e) GyroDeblurNet + curriculum learning 27.28 0.7803
Ablation study
We conduct an ablation study to investigate the effect of individual components in our method. Specifically, we first study the impact of using gyro data. To this end, we build a variant of GyroDeblurNet that has no gyro module, and uses only image features for computing deformable convolution offsets in the gyro deblurring blocks. In our method, we compensate for gyro errors using image features in the gyro refinement and gyro deblurring blocks. Thus, we also examine the effect of using image features in these blocks using two variants of GyroDeblurNet. Specifically, we build one variant without the connections from the image deblurring module to the gyro module, and another variant that computes deformable convolution offsets using only gyro features. We trained all the aforementioned models and our full model without curriculum learning. Finally, to study the impact of the curriculum-learning-based training scheme, we also include our full model trained with curriculum learning.
Tab. 4 reports the comparison result. As shown in Tab. 4 (a), deblurring with no gyro data performs the worst as it cannot benefit from the blur cue provided by gyro data. The results in Tab. 4 (b) and (c) show that exploiting gyro data enhances deblurring quality, but their performances are still worse than our full model in (d) as they do not fully compensate for gyro errors. Finally, the curriculum learning (Tab. 4 (e)) substantially improves the deblurring quality as it helps train the model to fully utilize gyro data.
6 Conclusion
In this paper, we proposed GyroDeblurNet, a novel gyro-based single image deblurring method that can effectively restore sharp images with the help of gyro data. To fully exploit gyro data while considering real-world gyro error, we presented a novel gyro refinement block, a novel gyro deblurring block, and a curriculum learning-based training strategy. In addition, we also presented the camera motion field, a novel gyro embedding scheme to represent real-world camera motions. Finally, we proposed synthetic and real datasets where blurred images are paired with gyro data. Extensive quantitative and qualitative experiments demonstrate the effectiveness of our approach.
Limitations and future work
Our method has a few limitations. While our method uses a fixed value for the hyperparameter $M$, the complexity of camera shakes may increase along with the exposure time, which means that a longer exposure time may require a larger value for $M$. Our method does not use the accelerometer, which is also equipped with most smartphones nowadays, and which may provide additional valuable information on the camera motion. Our image deblurring module adopts a relatively simple network architecture compared to recent non-gyro-based single image deblurring approaches, and this may limit the deblurring performance of our method. Extending our work to address these limitations would be an interesting future direction.

### v3

@@
 Abstract
-In this paper, we present GyroDeblurNet, a novel single image deblurring method that utilizes a gyro sensor to effectively resolve the ill-posedness of image deblurring.
-The gyro sensor provides valuable information about camera motion during exposure time that can significantly improve deblurring quality.
-However, effectively exploiting real-world gyro data is challenging due to significant errors from various sources including sensor noise, the disparity between the positions of a camera module and a gyro sensor, the absence of translational motion information, and moving objects whose motions cannot be captured by a gyro sensor.
-To handle gyro error, GyroDeblurNet is equipped with two novel neural network blocks: a gyro refinement block and a gyro deblurring block.
-The gyro refinement block refines the error-ridden gyro data using the blur information from the input image.
-On the other hand, the gyro deblurring block removes blur from the input image using the refined gyro data and further compensates for gyro error by leveraging the blur information from the input image.
+In this paper, we present GyroDeblurNet, a novel single-image deblurring method that utilizes a gyro sensor to resolve the ill-posedness of image deblurring.
+The gyro sensor provides valuable information about camera motion that can improve deblurring quality.
+However, exploiting real-world gyro data is challenging due to errors from various sources.
+To handle these errors, GyroDeblurNet is equipped with two novel neural network blocks: a gyro refinement block and a gyro deblurring block.
+The gyro refinement block refines the erroneous gyro data using the blur information from the input image.
+The gyro deblurring block removes blur from the input image using the refined gyro data and further compensates for gyro error by leveraging the blur information from the input image.
 For training a neural network with erroneous gyro data, we propose a training strategy based on the curriculum learning.
 We also introduce a novel gyro data embedding scheme to represent real-world intricate camera shakes.
-Finally, we present a synthetic dataset and a real dataset for the training and evaluation of gyro-based single image deblurring.
+Finally, we present both synthetic and real-world datasets for training and evaluating gyro-based single image deblurring.
 Our experiments demonstrate that our approach achieves state-of-the-art deblurring quality by effectively utilizing erroneous gyro data.
 1 Introduction
-Blur caused by camera shakes severely degrades image quality as well as the performance of various computer vision tasks such as classification, object detection, and semantic segmentation.
+Blur caused by camera shakes severely degrades image quality as well as the performance of various computer vision tasks.
 To overcome the image degradation caused by blur, single image deblurring has been extensively studied for decades [5, 11, 53, 33, 49, 16, 14].
-Recently, various DNN-based approaches have been proposed and have shown significant performance improvements over classical approaches [31, 53, 44, 23, 24, 51, 5, 4].
-Nonetheless, they still fail on images with large blur due to the severe ill-posedness of deblurring, i.e., multiple different blur trajectories could lead to similar-looking blurry images.
-To address the ill-posedness, several attempts have been made to exploit gyro sensors with which most smartphones are nowadays equipped.
-As gyro data are correlated with motion of the camera mounted on a smartphone, classical approaches based on blur models use gyro data to guide blur kernel estimation [17, 40, 41, 11, 14, 29].
-However, their performances are limited by their restrictive blur models that cannot effectively handle noise, nonlinear camera response functions, saturated pixels, moving objects, and so on.
-With the advance of deep learning, DNN-based approaches have also been proposed to exploit gyro sensors [30, 25, 16].
-Thanks to the remarkable capability of DNNs and gyro sensors, they show better performance compared to classical gyro-based approaches and recent DNN-based approaches without gyro sensors.
-However, recent DNN-based approaches that utilize gyro data are still limited in handling real-world blurred images.
-These approaches assume that the gyro data can accurately represent the blur in the blurred image, but this assumption does not hold in practice.
-First, gyro data from mobile devices such as smartphones usually have a significant amount of noise caused by various sources [19, 10, 17].
+Recently, various deep neural network (DNN)-based approaches have been proposed and have shown significant performance improvements over classical approaches.
+[31, 54, 44, 21, 22, 52, 6, 3].
+Nonetheless, they still fail on images with large blur due to the severe ill-posedness of deblurring.
+To address the ill-posedness of image deblurring, several attempts have been made to exploit the gyro sensors that most smartphones are now equipped with.
+Classical approaches based on blur models use gyro data to guide blur kernel estimation [16, 41, 42, 10, 14, 29].
+However, their performance is limited by restrictive blur models that struggle to handle noise, nonlinear camera response functions, saturated pixels, moving objects, and other challenges effectively.
+With the advancement of deep learning, DNN-based approaches have also been proposed to exploit gyro data [30, 23, 15, 34, 46].
+Thanks to the remarkable capabilities of DNNs and the rich information from gyro sensors, these methods demonstrate superior performance compared to the classical gyro-based approaches.
+Figure 1: Our method shows robust performance even with errors in real-world gyro data.
+(a) A real-world blurry image.
+(b) Gyro data visualization onto the blurry image.
+The gyro data do not perfectly match the blur trajectories due to errors.
+(c) Our deblurred result.
+However, recent DNN-based methods that utilize gyro data are still limited in handling real-world blurred images.
+First, gyro data from mobile devices such as smartphones usually contain a significant amount of noise caused by various sources [17, 8, 16].
 Second, gyro sensors measure the angular velocity around their center.
-Thus, to exploit gyro sensors, the recent gyro-based methods assume that the gyro sensor and the camera share the same center, and the camera is shaken only by rotational motions.
-However, in practice, the positions of the camera and the gyro sensor are different, and the camera shakes are caused by more complex motions.Third, real-world blurred images may have moving objects with different blur trajectories that cannot be captured by gyro sensors.
-All these factors make the motion information encoded in the gyro data inconsistent with the blur in the blurred image, and make it difficult to exploit the gyro data in real-world image deblurring.
-In this paper, we refer to such inconsistency between the gyro data and blur as gyro error.
-In this paper, we propose GyroDeblurNet, a novel gyro-based single image deblurring approach that can produce high-quality deblurring results in the presence of gyro error.
-To address the gyro error, GyroDeblurNet adopts a carefully designed network architecture that contains a novel gyro refinement block and a novel gyro deblurring block.
-The gyro refinement block refines gyro data with error using the blur information in the input image.
-On the other hand, the gyro deblurring block removes blur in the input image using the refined gyro data.
-To further compensate for gyro error, the gyro deblurring block also exploits the blur information in the input image as well as the input gyro data.
-Despite the proposed blocks, training a neural network with erroneous gyro data is difficult as the network could be trained to simply ignore the input gyro data.
-To resolve this issue, we also propose a curriculum-learning-based training strategy [2].
+Therefore, recent gyro-based methods assume that the gyro sensor and the camera share the same center and that the camera is only affected by rotational motions.
+However, in practice, the positions of the camera and the gyro sensor differ, and the camera experiences more complex shakes.
+Third, real-world blurred images may have moving objects with different blur trajectories that cannot be captured by gyro sensors.
+All these factors cause the motion information encoded in the gyro data to be inconsistent with the blur in the blurred image (Fig.
+1-(b)), making it challenging to exploit gyro data in real-world image deblurring.
+We refer to such inconsistency between the gyro data and blur as gyro error.
+In this paper, we propose GyroDeblurNet, a novel gyro-based single image deblurring approach that produces high-quality deblurring results even in the presence of gyro error (Fig.
+1-(c)).
+To address gyro error, GyroDeblurNet adopts a carefully designed network architecture that includes novel gyro refinement blocks and gyro deblurring blocks.
+The gyro refinement block corrects gyro data errors using the blur information from the input image.
+On the other hand, the gyro deblurring block removes blur from the input image using the refined gyro data.
+Additionally, to further compensate for gyro error, the gyro deblurring block leverages both the blur information in the input image and the input gyro data.
+Despite these blocks, training a neural network with erroneous gyro data is challenging, as the network might be trained to ignore the input gyro data and rely solely on the input image.
+To resolve this issue, we propose a curriculum-learning-based training strategy [1].
 To handle complex real-world camera shakes, we also present the camera motion field, a novel gyro data embedding scheme that can represent complex camera motions.
-Real-world blurry images often exhibit complex blur trajectories due to random camera shakes, which often occur when shooting a still shot with a long exposure time.
-Nonetheless, recent gyro-based approaches [30, 25, 16] assume smooth camera motions and represent camera motions using one or two motion vector per pixel or a couple of homographies for the entire image, which leads to low-quality deblurring results.
+Real-world blurry images often exhibit complex blur trajectories.
+Nonetheless, recent gyro-based approaches [30, 23, 15, 34] assume smooth camera motions and represent camera motions using one or two motion vectors per pixel or a couple of homographies for the entire image, which leads to low-quality deblurring results.
 To resolve this issue, our camera motion field is designed to represent complex camera motions using multiple vectors per pixel while maintaining a comparable memory footprint exploiting the spatially smooth nature of camera shakes.
-For the training and evaluation of gyro-based deblurring, previous methods [30, 25, 16] rely on synthetic datasets.
-However, these datasets do not accurately reflect the noise and blur in real-world blurred images as shown by Rim et al.
-[35].
-Additionally, their gyro data do not reflect real-world camera shakes that occur when capturing still-shot images either.
+Previous methods [30, 23, 15, 34] rely on synthetic datasets to train and evaluate the methods.
+However, their gyro data do not reflect real-world camera shakes that occur when capturing still-shot images either.
 Gyro data of previous methods were generated by random sampling [55] or obtained from the Visual-Inertial dataset [39].
-However, the Visual-Inertial dataset is originally developed for visual odometry and SLAM, so its camera motions are collected from constantly moving cameras, which differ from camera shakes of still-shot images.
-For training and evaluation of gyro-based deblurring for real-world blurred images, we propose two datasets: GyroBlur-Synth and GyroBlur-Real.
+However, the Visual-Inertial dataset was originally developed for visual odometry and SLAM, so its camera motions were collected from constantly moving cameras, which differ from camera shakes of still-shot images.
+For training and evaluating gyro-based deblurring for real-world blurred images, we propose two datasets: GyroBlur-Synth and GyroBlur-Real.
 GyroBlur-Synth is a synthetic dataset for both training and evaluation.
 The dataset consists of sharp and blurred image pairs with their corresponding gyro data.
-The gyro data are acquired from a smartphone device to reflect the real-world camera shakes.
-The blurred images are synthetically generated using the acquired gyro data and the blur synthesis process of RSBlur [34] for realistic blur.
-GyroBlur-Real is a real-world blur dataset for qualitative evaluation with no ground-truth sharp images.
-GyroBlur-Real provides 100 real-world blurred images paired with gyro data collected from a smartphone.
+The gyro data are acquired from a smartphone device to reflect real-world camera shakes.
+The blurred images are synthetically generated using the acquired gyro data and the blur synthesis process of RSBlur [36] for realistic blurred images.
+GyroBlur-Real is a real-world blur dataset for qualitative evaluation with no ground-truth sharp images, and it provides real-world blurred images paired with gyro data collected from a smartphone.
 We validate the performance of GyroDeblurNet on both synthetic and real-world datasets.
@@
 Our contributions can be summarized as follows:
-  • We propose a novel gyro-based single image deblurring approach, GyroDeblurNet.
-Extensive quantitative and qualitative evaluations show that GyroDeblurNet outperforms existing single image deblurring methods.
-  • To handle gyro error, we introduce a carefully designed network equipped with a novel gyro refinement block and a novel gyro deblurring block.
-We also develop a novel gyro data embedding scheme to represent complex camera shakes and propose a curriculum learning-based training scheme.
+  • We propose a novel gyro-based single image deblurring approach, GyroDeblurNet. Extensive quantitative and qualitative evaluations show that GyroDeblurNet outperforms existing deblurring methods.
+  • To handle gyro error, we introduce a carefully designed network equipped with novel gyro refinement and gyro deblurring blocks.
+We also develop a novel gyro data embedding to represent complex camera shakes and propose a curriculum learning-based training scheme.
   • We present synthetic and real image-gyro paired datasets with realistic camera motions, GyroBlur-Synth and GyroBlur-Real, for training and evaluating gyro-based single image deblurring methods.
@@
 Nevertheless, single image deblurring is still a challenging problem due to its high ill-posedness, and they still fail on images with large blur.
-To resolve the ill-posedness of deblurring, there have been attempts to leverage inertial measurement sensors such as the gyro sensor and accelerometer as they provide valuable information on the rotational and translational motion of a camera, respectively.
+To resolve the ill-posedness of deblurring, there have been attempts to leverage inertial measurement sensors such as the gyro sensor and accelerometer as they provide valuable information on the camera motion.
 Joshi et al.
@@
 They introduce a U-Net-based architecture [37] that takes a concatenation of a blurred image and the gyro data as input, which is also adopted by Lee et al.
-later [25].
+[23].
 Ji et al.
-[16] present a deformable-convolution-based network module for handling gyro data, which computes deformable convolution offsets from input gyro data to transform convolution filters according to the blur direction.
-However, these methods assume that the gyro data is accurate, neglecting potential errors.
-Moreover, they assume smooth camera motions and model camera motions using a motion vector field consisting of either one [30] or two motion vectors [16] per pixel, or a couple of homographies [25] representing the beginning and end position of the camera during the exposure time.
-Unfortunately, such restrictive assumptions do not hold for real-world blurred images with complex camera shakes as discussed in Sec.
+[15] adopt deformable convolutions to adjust convolution filters based on gyro data.
+Ren et al.
+[34] propose a transformer-based network that utilizes gyro data.
+However, these methods assume accurate gyro data, neglecting potential errors.
+Moreover, they assume smooth camera motions and model camera motions using a vector field consisting of either one [30, 34] or two vectors [15] per pixel, or a couple of homographies [23] representing the beginning and end position of the camera during the exposure time.
+Unfortunately, such restrictive assumptions do not hold in real-world cases as discussed in Sec.
 1.
-Utilizing gyro data by converting it into the form of motion vectors for single image deblurring can be considered as a non-blind deconvolution approach.
-While non-blind deconvolution methods have been studied for decades [33, 49, 22, 18, 26], they are not robust to blur kernel errors and produce unsatisfactory results with erroneous blur kernels.
-To overcome such limitation, several works adopted hand-crafted priors such as sparse prior [15], hyper-Laplacian prior [46, 22] and deep residual prior [43].
-However, they may fail to utilize erroneous blur kernels where such prior assumptions do not hold.
+Deblurring using gyro data can be framed as non-blind deconvolution, with gyro data serving as blur kernels.
+However, most non-blind deconvolution approaches assume accurate blur kernels without errors, making them unsuitable for erroneous gyro data [35, 50, 24, 20].
+Some attempts have been made to address kernel errors in non-blind deconvolution.
+Vasu et al.
+[47] apply varying regularization strengths in a non-blind deblurring algorithm and use a neural network to combine results and refine artifacts induced by kernel errors.
+Nan et al.
+[32] tackle kernel error artifacts by estimating residual errors induced by kernel errors using a CNN.
+However, these approaches assume mild errors in kernels and fail to handle the significant errors presented in gyro data, as will be demonstrated in our experiments.
+Figure 2: Camera motion field construction. (a) Computing homography and warped pixel coordinates. (b) Constructing camera motion field by stacking motion vectors.
 3 GyroDeblurNet
-The goal of GyroDeblurNet is to estimate a sharp deblurred image $S$ from an input blurred image $B$ and its corresponding gyro data $G$ possibly with large errors.
+The goal of GyroDeblurNet is to estimate a sharp deblurred image $D$ from an input blurred image $B$ and its corresponding gyro data $G$ possibly with large errors.
 Specifically, we define $G$ as a sequence of gyro data such that $G=\left\{g_{0},\cdots,g_{T-1}\right\}$, where $g_{t}$ is the $t$-th gyro data consisting of three angular velocities during the exposure time.
 $T$ is the number of gyro data, which is proportional to the exposure time.
-For example, $T=20$ if the exposure time is 0.1 seconds, and the sampling rate of the gyro sensor is 200 Hz.
 Instead of directly using $G$, we devise a novel gyro data embedding scheme named camera motion field to effectively handle complex camera shakes of an arbitrary exposure duration.
-In the rest of this section, we explain the gyro data embedding scheme based on the camera motion field, network architecture, and training strategy of GyroDeblurNet in detail.
-Figure 1: Camera motion field construction. (a) Computing homography and warped pixel coordinates. (b) Constructing camera motion field by stacking motion vectors.
+In the rest of this section, we explain the gyro data embedding scheme, network architecture, and training strategy of GyroDeblurNet in detail.
 Figure 3: Network architecture of GyroDeblurNet.
-Figure 3: Detailed architectures of the gyro refinement block and gyro deblurring block.
+Figure 4: Detailed architectures of the modules in the GyroDeblurNet. (a) Architecture of the gyro refinement block. (b) Architecture of the gyro deblurring block.
 3.1 Gyro Data Embedding
-GyroDeblurNet first converts input gyro data sequence $G$ into a camera motion field $\mathcal{V}$ before feeding it into the subsequent neural network to handle temporally complex camera shakes of an arbitrary length $T$.
+GyroDeblurNet first converts input gyro data sequence $G$ into a camera motion field $\mathcal{V}$ to handle temporally complex camera shakes of an arbitrary length $T$.
 The conversion process is illustrated in Fig.
-1.
-Our camera motion field has a fixed channel size $2M$ where $M$ is a hyperparameter, regardless of the length of the input gyro data to allow it to be fed to convolutional neural networks.
+2.
+A camera motion field has a fixed channel size $2M$ where $M$ is a hyperparameter representing the number of vectors in camera motion field, regardless of the length of the input gyro data to allow it to be fed to convolutional neural networks.
 To achieve this, we first resample the input gyro data sequence $G$ of an arbitrary length $T$ to $M+1$ samples using cubic-spline interpolation.
 We denote the resampled sequence as $G^{\prime}=\left\{g^{\prime}_{0},\cdots,g^{\prime}_{M}\right\}$ where $g^{\prime}_{m}$ is a resampled gyro sample.
-The hyperparameter $M$ is set to an even number so that we have the same number of gyro data before and after the temporal center of the exposure duration.
+The hyperparameter $M$ is set to an even number so that we have the same number of gyro data before and after the temporal center.
 Then, assuming that the camera and the gyro shares the same center and there exists neither off-center rotational nor translational motions, we integrate the angular velocities and obtain $M+1$ camera orientations $\theta_{m}$, where $m=0,\cdots,M$, including the camera orientations at the beginning and end of the exposure.
@@
 For each $\theta_{m}$, we compute a homography $H_{m}$ as $H_{m}=KR(\theta_{m})K^{-1}$ where $K$ is the camera intrinsic matrix and $R(\theta_{m})$ is the rotation matrix corresponding to $\theta_{m}$.
-Based on the camera orientation that we chose as a reference, there can be multiple sets of homography $H_{m}$ each of which leads to shifted vector field.
-To prevent unnecessary shift after deblurring, we assume $\theta_{M/2}$ at the temporal center to be $(0,0,0)$ so that $H_{M/2}$ is an identity matrix when computing $\theta_{m}$.
+To prevent unnecessary shift after deblurring, we chose the temporal center as a reference by assuming $\theta_{M/2}$ at the temporal center to be $(0,0,0)$ so that $H_{M/2}$ is an identity matrix when computing $\theta_{m}$.
 Finally, from the obtained homographies, we compute a camera motion field $\mathcal{V}$, which illustrates how each pixel is blurred.
 Specifically, we define a camera motion field $\mathcal{V}$ as a tensor of size $W/{s}\times H/{s}\times 2M$ where $W$ and $H$ are the width and height of the input blurred image, and $s$ is a scaling factor.
-Then, at each spatial position $(x,y)$ in $\mathcal{V}$, we compute the warped coordinates of $(x,y)$ by $H_{m}$ for all $m$ and compute the difference between temporally consecutive coordinates to obtain $M$ motion guidance vectors at each position.
+Then, at each position $(x,y)$ in $\mathcal{V}$, we compute the warped coordinates of $(x,y)$ by $H_{m}$ for all $m$ and compute the difference between temporally consecutive coordinates to obtain $M$ vectors.
 By stacking the vectors for all the spatial positions, we construct $\mathcal{V}$.
 In our experiments, we set $M=8$ and $s=2$.
-Our camera motion field is motivated by the blur field proposed by Mustaniemi et al.
-[30] and the motion guidance vector (MGV) field introduced by Ji et al.
-[16].
-Unlike previous methods that rely on only one or two vectors to represent camera shakes at each pixel, our camera motion field offers a more sophisticated representation.
+Fig.
+1-(b) visualizes camera motion field vectors onto a real-world blurry image.
+Note that similar representations for gyro data based on vector fields have also been proposed by previous approaches [30, 15], as discussed in Sec.
+2.
+However, unlike these methods that rely on only one or two vectors to represent camera shakes at each pixel, our camera motion field offers a more sophisticated representation.
 By adjusting the hyperparameter $M$, it can effectively capture temporally intricate camera shake patterns.
@@
 3.2 Network Architecture
-The network architecture of GyroDeblurNet is illustrated in Fig.
-2.
-GyroDeblurNet takes a blurred image $B$ and a camera motion field $\mathcal{V}$ as input and estimates a deblurred image $S$.
-The network consists of two modules: an image deblurring module and a gyro module.
-The image deblurring module takes a blurred image $B$ and performs deblurring with the aid of gyro data.
-It then produces a residual $R$ that is added back to $B$ to produce $S$.
+GyroDeblurNet takes a blurred image $B$ and a camera motion field $\mathcal{V}$ as input and estimates a deblurred image $D$.
+The network consists of two modules: an image deblurring module and a gyro module (Fig.
+3).
+The image deblurring module deblurs a blurred image $B$ with the aid of gyro data.
+It then produces a residual $R$ that is added back to $B$ to produce $D$.
 The image deblurring module adopts a U-Net architecture [37] and the NAFBlock [3] as its basic building block.
-The image deblurring module adopts gyro deblurring blocks in its bottleneck to perform deblurring using gyro features from the gyro module.
+The image deblurring module adopts gyro deblurring blocks in its bottleneck to perform deblurring using gyro features from the gyro network.
 On the other hand, the gyro module takes a camera motion field $\mathcal{V}$ and refines it with the aid of image features from the image deblurring module.
-The gyro module consists of a convolution layer to embed an input camera motion field $\mathcal{V}$ into the feature space, and gyro refinement blocks.
+The gyro module consists of a convolution layer to embed an input camera motion field $\mathcal{V}$ into the feature space, gyro refinement blocks and strided convolution layers.
 The gyro refinement blocks refine gyro features with the aid of image features from the encoder of the image deblurring module.
@@
 For a detailed network architecture including feature dimensions, refer to the supplementary material.
-Figure 4: The input camera motion field passes a convolutional layer to produce differently refined camera motion candidates in different channels.
-The candidates are then weighted by computing similarity with the blur information of the blurred image.
 Gyro refinement block
-The structure of the gyro refinement block is illustrated in Fig.
-3.
+Fig.
+4-(a) illustrates the structure of the gyro refinement block.
+The camera motion information in the input camera motion field $\mathcal{V}$ may contain significant global errors caused by gyro sensor noise, unknown rotational centers, and missing translational motions, as visualized in Fig.
+1-(b).
+The gyro refinement block is designed to compensate for these global errors using the blur information in the input image based on the following intuition.
 The gyro refinement block takes a gyro feature computed from the input camera motion field $\mathcal{V}$ as input.
-Then, it refines the gyro feature using the channel attention mechanism [13] with the aid of an image feature from the encoder of the image deblurring module.
-Specifically, the gyro refinement block takes an image feature from the image deblurring module as an additional input, and concatenates it with the input gyro feature.
-It then computes channel-wise attention weights through global average pooling and $1\times 1$ convolution layers, and refines the gyro feature using the obtained channel-wise attention weights.
-Finally, it applies a convolution layer to further refine the gyro feature.
-The intuition behind the refinement process by the gyro refinement block is illustrated in Fig.
-4.
-The camera motion information in the input camera motion field $\mathcal{V}$ may contain a significant amount of error.
-From such an erroneous camera motion information, the convolution layer at the beginning of the gyro module produces a gyro feature consisting of multiple camera motion candidates at different channels by introducing different perturbations to the input camera motion information.
-Then, the gyro refinement block adaptively selects the channels of the gyro feature that are consistent with the blur information encoded in the image feature using the channel attention weights computed from both gyro and image features [13].
+This input gyro feature encodes multiple camera motion candidates across different channels, obtained by introducing perturbations to the input camera motion information in the previous layers.
+To globally refine the gyro feature, the gyro refinement block adaptively selects the channels of the gyro feature that are consistent with the blur information in the input image.
+Specifically, the gyro refinement block takes an image feature from the image deblurring module as an additional input and concatenates it with the input gyro feature.
+It then computes channel-wise weights [13] through global average pooling and $1\times 1$ convolution layers, and multiplies them to the gyro feature.
+Following the gyro refinement block, a strided convolution layer further refines the gyro feature while reducing its spatial resolution.
 Gyro deblurring block
-The structure of the gyro deblurring block is shown in Fig.
-3.
-The gyro deblurring block performs deblurring of the image feature with the gyro feature from the gyro module.
+The gyro deblurring block, illustrated in Fig.
+4-(b), performs deblurring of the image feature using the gyro feature from the gyro module.
 The input image $B$ may have spatially-variant blur caused by camera shakes and moving objects.
-To effectively remove such spatially-varying blur in the image feature, the gyro deblurring block adopts the deformable convolution operation [7, 55], whose offsets are computed based on both image and gyro features.
-As the gyro feature provides accurately refined information on the camera shakes, and the image feature provides information on the blur in $B$ including object motion blur, the gyro deblurring block can successfully handle both camera shakes, and object motion blur.
-To further remove the remaining spatially-variant blur in the deblurred image feature, the gyro deblurring block applies a spatial attention block to the deblurred image feature following Xu et al.
-[50].
-The deblurred feature is then further refined through a NAFBlock and a convolution layer.
-While deformable convolution has also been employed by EggNet [16] to deblur images using gyro data, EggNet applies it without accounting for real-world gyro errors.
-This oversight leads to subpar results for real-world images, as will be shown in Sec.
-5.
-In contrast, our approach incorporates network modules, a training strategy, and data specifically designed to accommodate real-world gyro errors.
+To effectively remove such blur, the gyro deblurring block consists of two sub-blocks: the first sub-block performs spatial refinement of the blur information and deblurring of the image feature using the refined blur information, while the second sub-block further refines the deblurred image feature.
+Specifically, the first sub-block concatenates the image feature and gyro feature, computing spatially-refined blur information in the form of deformable convolution kernel offsets.
+It then performs deblurring of the image feature using these offsets.
+The second sub-block further refines the deblurred image feature using spatial attentions [51].
+Spatial attention weights are computed from the deblurred image block through convolution and sigmoid layers, and multiplied with the deblurred image feature, which is further processed through convolution and concatenation layers.
 3.3 Training Strategy
-Despite the gyro refinement block and gyro deblurring block, it is still challenging to train a deblurring network with erroneous gyro data.
-Naïvely training the network with erroneous gyro data can lead the network to ignore the gyro data and simply rely only on the blur information in the blurred image.
-Thus, to train our network to fully utilize input gyro data, we introduce a training strategy based on curriculum learning [2], which has shown superior performance in various learning-based approaches that deal with erroneous data [37, 39, 42].
+To address the challenge of training a deblurring network with erroneous gyro data, we propose a curriculum learning-based training strategy.
+This approach, inspired by its success in handling noisy data in prior work [1, 38, 40, 43], aims to progressively guide the network in leveraging gyro information effectively despite inaccuracies.
 Our training strategy initially trains the network with error-free gyro data, and gradually increases error in the training data.
-To this end, we first assume that for each blurred image $B$, its corresponding error-free gyro data $G_{GT}$ is given as well as its ground-truth sharp image $S_{GT}$.
+To this end, we first assume that, for each blurred image $B$, its corresponding error-free gyro data $G_{GT}$ is given as well as its ground-truth sharp image $S_{GT}$.
 From $G_{GT}$, we compute a noise-free camera motion field $\mathcal{V}_{clean}$ and a noisy camera motion field $\mathcal{V}_{noisy}$.
@@
 We then compute a blended camera motion field $\mathcal{V}_{\alpha}=(1-\alpha)\cdot\mathcal{V}_{clean}+\alpha\cdot\mathcal{V}_{noisy}$ where $\alpha$ is a blending parameter.
-Finally, we train our network with a PSNR loss [4, 3], which is defined as:
-$\mathcal{L}_{PSNR}=10\cdot\log_{10}(\|f(B,\mathcal{V}_{\alpha})-S_{GT}\|^{2}+\epsilon)$ (1)
-where $f(B,\mathcal{V}_{\alpha})$ represents a deblurred output of GyroDeblurNet and $\epsilon$ is a small constant to prevent the negative infinity, which is set to $81e-$ 8 in our experiments.
-We initially set $\alpha$ to 0, and gradually increase it to $1$ during training.
+We gradually increase $\alpha$ from 0 to 1 during training.
 For more details including the noisy camera motion field generation and the scheduling of $\alpha$, refer to the supplementary material.
 4 GyroBlur Dataset
-Figure 5: Examples of the GyroBlur-Synth and GyroBlur-Real datasets. In (a), the top and bottom rows show ground-truth images and the corresponding blurry images, respectively.
 4.1 GyroBlur-Synth
 GyroBlur-Synth provides 14,600 and 640 synthetically blurred images of size $720\times 1280$ for the training and test sets.
-The dataset also provides ground-truth sharp images, ground-truth clean gyro data.
-To faithfully mimic real-world blurred images, the dataset covers various sources of image degradation, such as camera shakes, moving objects, noise, and saturation.
-The dataset also includes gyro error data to produce realistic erroneous gyro data from the clean gyro data.
-Fig.
-5 (a) shows examples of GyroBlur-Synth.
+The dataset also provides ground-truth sharp images and the corresponding gyro data.
+To faithfully mimic real-world blurred images, the dataset covers various degradation sources, such as moving objects, noise, and saturation.
 To build a dataset with realistic camera motions, we implemented an Android application that records gyro data on a mobile phone.
@@
 We then add the object to the warped sharp images before blending them during the blur synthesis process described earlier.
-This method may seem simple, but our experiment shows that it helps train our model to better handle spatially-variant motion blur from moving objects.
 To reflect real-world noise and saturated pixels, we also adopt the blur synthesis pipeline of RSBlur [36].
@@
 Then, we synthesize saturated pixels, convert the image into the camera RAW space, synthesize noise, and convert the noisy RAW image to the linear sRGB image.
+We use the camera parameters of a Samsung Galaxy S22 ultra-wide camera, and the noise distribution estimated from it.
 We refer the readers to [36] for the detailed blur synthesis process.
-The blur synthesis pipeline of RSBlur requires camera parameters such as a color correction matrix and a noise distribution.
-To this end, we use the camera parameters of a Samsung Galaxy S22 ultra-wide camera, and the noise distribution estimated from it.
-The gyro error data includes gyro noise and randomly perturbed rotation centers for generating noisy camera motion fields.
+To generate noisy camera motion fields, we add gyro errors with gyro data noise and randomly perturbed rotation centers.
 To obtain realistic gyro noise, we measured the gyro noise distribution from a stationary smartphone (Samsung Galaxy S22) placed on a table.
@@
 For more details on the dataset, we refer the readers to the supplementary material.
-Our synthetic dataset generation approach requires minimal device-specific information, including a small amount of gyro data sequences and parameters.
-This approach minimizes the need for laborious manual data collection.
-Specifically, we can collect a sufficient amount of gyro data by holding a smartphone in hand only for a couple of minutes, and by placing a smartphone on a table only for a few seconds.
-The camera parameters required for the RSBlur pipeline can also be easily obtained by a simple calibration process with a couple of images captured by a target camera.
-Note also that our synthetic dataset approach process does not explicitly model the relative position of the gyro sensor with respect to the camera module, which is information that is hard to access.
-Instead, our approach implicitly models it as part of rotational center error.
-Despite this, our experiments show that our approach can effectively remove blur.
+Our synthetic dataset generation requires minimal device-specific input, brief gyro sequences, and simple camera calibration, significantly reducing manual data collection.
+We capture sufficient gyro data within minutes and obtain the necessary camera parameters for the RSBlur pipeline from a few calibration images.
+While we do not explicitly model the gyro’s position relative to the camera, this is implicitly addressed as gyro errors such as rotational center error.
+Our experiments demonstrate effective blur removal regardless of these errors.
+Method Models GyroBlur-Synth GyroBlur-Real-S Param. (M) MACs (G) Time (s)
+PSNR $\uparrow$ SSIM $\uparrow$ NIQE $\downarrow$ TOPIQ $\uparrow$
+Single-image deblurring MPRNet [52] 25.03 0.7081 5.25 0.397 20.13 10927 0.998
+NAFNet [4] 25.06 0.7085 5.27 0.409 17.11 227 0.106
+AdaRevD-B (NAFNet) [27] 25.61 0.7281 5.18 0.425 43.81 2658 1.661
+Uformer-B [48] 25.72 0.7334 4.80 0.431 50.88 2143 1.061
+Stripformer [45] 25.93 0.7398 4.71 0.456 19.71 2397 1.367
+FFTformer [19] 26.01 0.7481 4.98 0.434 16.56 1894 1.879
+Non-blind Vasu et al. [47] 19.63 0.4526 6.13 0.317 10.41 - 160.176
+deblurring Nan et al. [32] 22.22 0.5313 5.81 0.348 26.16 41170 27.069
+Gyro-based deblurring DeepGyro [30] 23.78 0.6649 5.64 0.381 31.03 769 0.068
+EggNet [15] 25.49 0.7266 5.18 0.413 6.34 1102 0.071
+INformer [34] 25.11 0.7103 5.29 0.408 24.88 1315 0.464
+Ours 27.28 0.7803 4.47 0.548 16.31 262 0.130
+Table 1: Quantitative comparison on the GyroBlur-Synth and GyroBlur-Real-S.
+For the GyroBlur-Real-S, we use non-reference metrics (NIQE [28] and TOPIQ [2] trained on KonIQ-10k dataset [12]) for evaluation.
+Inference times were measured using a $720\times 1280$ image.
+For models utilizing gyro data, inference times include both gyro data embedding construction and the deblurring process.
+Figure 5: Qualitative comparison on GyroBlur-Synth. The red curves overlaid on the blurred images visualize the input gyro data.
 4.2 GyroBlur-Real
-GyroBlur-Real provides 100 real-world blurry images and their corresponding gyro data for qualitative evaluation.
-The dataset was collected using a Samsung Galaxy S22.
-The blurry images are of size $4080\times 3060$, and provided in both raw DNG and JPEG formats.
-The dataset also provides metadata such as timestamps of the images and gyro data, camera intrinsic parameters, exposure times and ISO parameters.
-Fig.
-5 (b) shows examples of GyroBlur-Real.
+GyroBlur-Real consists of two subsets: GyroBlur-Real-S and GyroBlur-Real-H.
+GyroBlur-Real-S was collected using a Samsung Galaxy S22, which was also used for constructing the GyroBlur-Synth dataset.
+We use GyroBlur-Real-S as the primary dataset to evaluate the performance of our approach on real-world images.
+On the other hand, GyroBlur-Real-H was collected using a Huawei P30 Pro, and is utilized for evaluating the generalization ability of our approach.
+GyroBlur-Real-S provides 100 real-world blurry images of size $4080\times 3060$ along with their corresponding gyro data.
+GyroBlur-Real-H provides 17 real-world blurry images of size $5120\times 3840$ and their gyro data.
+The images in both datasets are provided in DNG and JPEG.
 5 Experiments
-For evaluation, we implemented our network using PyTorch.
+We implemented our network using PyTorch.
 We trained our network on randomly cropped $256\times 256$ image patches from the GyroBlur-Synth dataset for 300 epochs with a batch size of 16.
@@
 The learning rate was initially set to 0.0001 and reduced to 1e-7 using the cosine annealing scheduler [26].
+We used the PSNR loss [3, 4] to train our model.
 We measured the computation times of all the models on a GeForce RTX 3090 GPU.
-One well-known issue in using gyro data is that the timestamps of the gyro sensor and the camera are not synchronized.
-Fortunately, several approaches that can successfully synchronize the gyro sensor and camera have been proposed [9].
-Thus, in our work, we assume that the input gyro data are already synchronized for the exposure time of the input image.
-GyroBlur-Real dataset was collected using an open-source Android application [1] that supports synchronization between the gyro and camera based on gyro-based clock synchronization [9].
-In the remainder of this section, we present the performance evaluation of GyroDeblurNet.
 We evaluate the performance of our method using the test set of GyroBlur-Synth and GyroBlur-Real.
 Additional experiments and examples including visualization of the effect of gyro refinement and an analysis on the robustness to gyro errors are included in the supplementary material.
-Table 1: Quantitative comparison on the test of GyroBlur-Synth.
-The inference times of the models that use gyro data include both gyro data embedding (e.g., camera motion field construction) and deblurring processes.
-  Gyro Models PSNR SSIM Param. (M) MACs (G) Time (s)
-X MPRNet [51] 25.03 0.7081 20.13 10927 0.998
-NAFNet [3] 25.06 0.7085 17.11 227 0.106
-Uformer-B [47] 25.72 0.7334 50.88 2143 1.061
-Stripformer [45] 25.93 0.7398 19.71 2397 1.367
-FFTformer [21] 26.01 0.7481 16.56 1894 1.879
-O DeepGyro [30] 23.78 0.6649 31.03 769 0.068
-EggNet [16] 25.49 0.7266 6.34 1102 0.071
-Ours 27.28 0.7803 16.31 262 0.130
-Figure 6: Qualitative comparison on GyroBlur-Synth. The red curves overlaid on the blurred images visualize the input gyro data.
-Figure 7: Qualitative comparison on GyroBlur-Real. The red curves overlaid on the blurred images visualize the input gyro data.
+Figure 6: Qualitative comparison on GyroBlur-Real-S. The red curves overlaid on the blurred images visualize the input gyro data.
 Comparison with state-of-the-art methods
-We first compare GyroDeblurNet with state-of-the-art single image deblurring methods that use gyro data  [30, 16] and those that do not [51, 3, 47, 45, 21].
-We evaluate the performance of all the models on the test set of GyroBlur-Synth with gyro errors.
+We compare GyroDeblurNet with state-of-the-art single image deblurring methods that use gyro data [30, 15, 34] and those that do not [52, 4, 48, 45, 19, 27].
+We also compare GyroDeblurNet with non-blind deconvolution methods that are designed to handle kernel error [32, 47].
+Since these methods assume uniform blur kernels, we extended them to handle spatially-varying blur kernels computed from gyro data by using patch-wise blur kernels [9].
+For more details about our extensions of the non-blind deconvolution methods, we refer the readers to the supplementary material.
+We train all models with the training set of GyroBlur-Synth and evaluate the performance of all the models on the test set of GyroBlur-Synth and GyroBlur-Real-S.
 The results are presented in Tab.
 1.
-All the models in the table were trained with GyroBlur-Synth using their own hyperparameter settings.
-As shown in Tab.
-1, GyroDeblurNet outperforms all the previous non-gyro-based methods by a large margin.
-Especially compared to Uformer-B [47], Stripformer [45] and FFTformer [21], which adopt transformers, our method requires a significantly smaller amount of computation, and a shorter computation time, while achieving more than 1 dB higher PSNR.
-Moreover, our method outperforms all the previous gyro-based methods by a large margin, while ours requires a comparable model size and a much smaller computational overhead.
-Qualitative comparisons on GyroBlur-Synth and GyroBlur-Real are shown in Fig.
+As shown in the table, GyroDeblurNet significantly outperforms all previous non-gyro-based methods.
+Specifically, compared to transformer-based approaches [48, 45, 19], our method requires substantially less computation and shorter computation time, while achieving more than 1 dB higher PSNR.
+Furthermore, our method surpasses all previous gyro-based methods by a large margin, despite requiring a comparable model size and much smaller computational overhead.
+The non-blind deblurring methods designed to handle blur kernel errors also perform significantly worse than our method due to large gyro errors.
+Nonetheless, our approach demonstrates superior performance even in the presence of large gyro errors.
+Qualitative comparisons on GyroBlur-Synth and GyroBlur-Real-S are shown in Fig.
 5 and in Fig.
-7, respectively.
+6, respectively.
 As shown in the figures, the previous non-gyro-based methods fail to restore sharp details due to severe blur.
 Moreover, despite the input gyro data, the previous gyro-based methods also fail to produce sharp details as they are not robust to gyro errors.
-On the other hand, our method successfully restores sharp images as our method can effectively exploit gyro data with errors.
-Figure 8: Camera motion field visualization on a GyroBlur-Synth image and a GyroBlur-Real image.
-(a) camera motion field visualization on a GyroBlur-Synth image.
-(b) Our result of (a).
-(c) A GyroBlur-Real image.
-(d) Gyro errors.
-(e) Our result of (c).
-Gyro errors
-As discussed in Sec.
-1, real-world gyro data cannot accurately represent the camera motion due to gyro noise and missing information such as rotational centers.
+The non-blind deblurring methods that are designed to handle kernel errors show significant ringing artifacts due to large errors in the gyro data.
+On the other hand, our method successfully restores sharp images as our method can effectively exploit erroneous gyro data.
+Model GyroBlur-Synth GyroBlur-Real-S
+PSNR $\uparrow$ / SSIM $\uparrow$ NIQE $\downarrow$ / TOPIQ $\uparrow$
+(a) Deblurring w/ no gyro data 24.90 / 0.6997 5.23 / 0.391
+(b) Training w/ error-free gyro data 24.94 / 0.7112 5.27 / 0.396
+(c) Deblurring w/ gyro data w/o any refine. 25.47 / 0.7128 5.01 / 0.418
+(d) Gyro refine. w/o image features 26.17 / 0.7465 4.75 / 0.447
+(e) Deform. conv. using only gyro features 26.32 / 0.7542 4.76 / 0.460
+(f) GyroDeblurNet 26.94 / 0.7667 4.61 / 0.511
+(g) GyroDeblurNet + curriculum learning 27.28 / 0.7803 4.47 / 0.548
+Table 2: Ablation study on the components of GyroDeblurNet. For the GyroBlur-Real-S, we use non-reference metrics (NIQE [28] and TOPIQ [2] trained on KonIQ-10K dataset [12]) for evaluation.
+Figure 7: Qualitative results of the ablation study on a real-world blurry image.
+Ablation study
+We conducted an ablation study to investigate the effects of individual components in our method.
+To this end, we built several variants of GyroDeblurNet and compared their performances.
+Both quantitative and qualitative results are reported in Tab.
+2 and Fig.
+7, respectively.
+Tab.
+2-(a) corresponds to a model with no gyro modules, using only image features in the gyro deblurring blocks.
+Tab.
+2-(b) represents our full model, but trained with error-free gyro data.
+Tab.
+2-(c) is a model that is trained with erroneous gyro data but does not refine gyro features using image features.
+Specifically, in this model, gyro features are not concatenated with image features in both gyro refinement and gyro deblurring blocks.
+Tab.
+2-(f) is our full model trained without curriculum learning, while Tab.
+2-(g) is our full model trained with curriculum learning.
+As shown in Tab.
+2-(a), deblurring with no gyro data performs the worst, as it cannot benefit from the blur cue provided by gyro data.
+The model in (b) achieves only marginal improvement over (a), indicating the importance of reflecting gyro errors in training data.
+The results in (c), (d), (e) and (f) demonstrate the necessity of training with erroneous gyro data and gyro refinement using image features.
+Finally, curriculum learning in (g) further improves deblurring quality, as it helps the model to fully utilize erroneous gyro data.
 Fig.
-8 shows examples of such gyro errors in our datasets where camera motion fields are visualized as red curves.
-As shown in the figure, the camera motion information in the camera motion fields in both GyroBlur-Synth and GyroBlur-Real do not perfectly match the blur trajectories in the blurred images due to gyro errors.
-Nonetheless, GyroDeblurNet successfully restores sharp images for both synthetic and real-world input images.
-Table 2: Results on GyroBlur-Synth with two models trained with noisy and noise-free gyro data.
-    Gyro data Noisy Noise-free
-  PSNR 27.28 24.94
-  SSIM 0.7803 0.7112
-Effect of training with noisy gyro data
-Our training data includes gyro errors to generalize our model to real-world gyro errors.
-Here we demonstrate the effect of the synthetically added gyro errors.
+7 also highlights the impact of each component in our method.
+The results in Fig.
+7-(b) and (c) reveal that training with error-free gyro data yields no performance gain over training without gyro data since error-free gyro data cannot reflect real-world gyro errors.
+In contrast, Fig.
+7-(d) shows that training with erroneous gyro data aids in handling these errors but remains suboptimal.
+Gyro refinement using image features in Fig.
+7-(e) significantly improves image quality, while Fig.
+7-(f) demonstrates how curriculum learning further enhances error handling.
+Figure 8: Results of a real-world blurred image with a moving object. (a) Blurry image and its background region and foreground object. (b) Result of FFTformer. (c) Result of GyroDeblurNet.
+Moving objects
+Another source of gyro error is moving objects with different blur trajectories.
 Fig.
-9 shows that the model trained with noisy gyro data successfully removes blur from a real-world blurred image with gyro errors, while the one trained with noise-free gyro data fails.
-Tab.
-2 presents a quantitative comparison on GyroBlur-Synth.
-As shown in the table, training with noisy gyro data substantially improves the deblurring quality by more than 2 dB, proving that including gyro errors in the training data is crucial for the training of gyro-based deblurring.
-Figure 9: Effect of training with noisy gyro data. (a) Real-world blurred image. (b) Model trained with noisy gyro data. (c) Model trained with noise-free gyro data.
-Moving objects
-Another source of gyro error is moving objects with different blur trajectories as discussed in Sec.
-1.
-Fig.
-10 shows a real-world example where the input blurry image has a moving object with different blur trajectory.
-Due to severe blur, the result of FFTformer [21], the state-of-the-art non-gyro-based method, still has remaining blur in both background and moving object.
-EggNet [16] also fails to deblur the image as it does not consider neither moving objects nor gyro errors.
-In contrast, our method successfully restores sharp details both on the moving object and the background thanks to its robustness to gyro errors.
-Figure 10: Results of a real-world blurred image with a moving object.
-Table 3: Effect of camera motion field hyperparameter $M$.
+8 shows a real-world example where the input blurry image has a moving object with a different blur trajectory.
+Due to severe blur, the result of FFTformer [19], the state-of-the-art non-gyro-based method, still has remaining blur in the background.
+In contrast, our method successfully restores sharp details on the background and shows comparable quality on the moving object thanks to its robustness to gyro errors.
+Figure 9: Effect of increasing $M$ for real-world complex motion blur. (a) Real-world blurry image with complex blur. (b) Result with $M=2$. (c) Result with $M=8$.
   $M$ 2 4 8 16
@@
 SSIM 0.7706 0.7651 0.7803 0.7811
+Table 3: Effect of camera motion field hyperparameter $M$.
 Camera motion field hyperparameter $M$
 We investigate the impact of the hyperparameter $M$ that determines the temporal resolution of the camera motion field.
-To verify the effect of $M$, we compare the deblurring performance of different values of $M$.
 As shown in Tab.
@@
 However, increasing $M$ to 16 does not yield a significant performance gain while it doubles the memory consumption, indicating that $M=8$ is sufficient to represent the camera shakes in most of the images in the test set of GyroBlur-Synth.
-Figure 11: Effect of increasing $M$ for real-world complex motion blur. (a) Real-world blurry image with complex blur. (b) Result with $M=2$. (c) Result with $M=8$.
-Table 4: Ablation study on the components of GyroDeblurNet.
-  Model PSNR SSIM
-(a) Deblurring with no gyro data 24.90 0.6997
-(b) Gyro refine. w/o image features 26.64 0.7584
-(c) Deform. conv. using only gyro features 26.68 0.7588
-(d) GyroDeblurNet 26.94 0.7667
-(e) GyroDeblurNet + curriculum learning 27.28 0.7803
-Ablation study
-We conduct an ablation study to investigate the effect of individual components in our method.
-Specifically, we first study the impact of using gyro data.
-To this end, we build a variant of GyroDeblurNet that has no gyro module, and uses only image features for computing deformable convolution offsets in the gyro deblurring blocks.
-In our method, we compensate for gyro errors using image features in the gyro refinement and gyro deblurring blocks.
-Thus, we also examine the effect of using image features in these blocks using two variants of GyroDeblurNet.
-Specifically, we build one variant without the connections from the image deblurring module to the gyro module, and another variant that computes deformable convolution offsets using only gyro features.
-We trained all the aforementioned models and our full model without curriculum learning.
-Finally, to study the impact of the curriculum-learning-based training scheme, we also include our full model trained with curriculum learning.
-Tab.
-4 reports the comparison result.
-As shown in Tab.
-4 (a), deblurring with no gyro data performs the worst as it cannot benefit from the blur cue provided by gyro data.
+Generalization to other devices
+We test GyroDeblurNet’s generalization ability by comparing it with top deblurring methods from Tab.
+1 on GyroBlur-Real-H.
+All methods were trained on GyroBlur-Synth.
 The results in Tab.
-4 (b) and (c) show that exploiting gyro data enhances deblurring quality, but their performances are still worse than our full model in (d) as they do not fully compensate for gyro errors.
-Finally, the curriculum learning (Tab.
-4 (e)) substantially improves the deblurring quality as it helps train the model to fully utilize gyro data.
+4 show our approach outperforms the others, even on a different device.
+Fig.
+10 also demonstrates that our approach successfully removes blur from an image captured by a different device, highlighting its generalization ability.
+Figure 10: Blurry image taken with a Huawei P30 Pro device and its deblurred result. (a) A blurry image and its gyro errors. (b) Deblurred result of (a).
+  Model Stripformer [45] FFTformer [19] EggNet [15] Ours
+NIQE $\downarrow$ 5.47 5.22 5.71 4.94
+TOPIQ $\uparrow$ 0.451 0.459 0.432 0.521
+Table 4: Quantitative evaluation on GyroBlur-Real-H using non-reference metrics [28, 2].
 6 Conclusion
@@
 While our method uses a fixed value for the hyperparameter $M$, the complexity of camera shakes may increase along with the exposure time, which means that a longer exposure time may require a larger value for $M$.
-Our method does not use the accelerometer, which is also equipped with most smartphones nowadays, and which may provide additional valuable information on the camera motion.
+Our method does not use the accelerometer, which is also equipped with most smartphones, and which may provide additional valuable information on the camera motion.
 Our image deblurring module adopts a relatively simple network architecture compared to recent non-gyro-based single image deblurring approaches, and this may limit the deblurring performance of our method.
 Extending our work to address these limitations would be an interesting future direction.
+Acknowledgements
+This work was supported by the National Research Foundation of Korea (NRF) grants (RS-2023-NR077065, RS-2023-00280400, RS-2023-00211658, RS-2024-00438532) and by Institute of Information & communications Technology Planning & Evaluation (IITP) grant (RS-2019-II191906, Artificial Intelligence Graduate School Program (POSTECH)) funded by the Korea government (MSIT).
+This work was also supported by Samsung Research Funding & Incubation Center of Samsung Electronics under Project Number SRFC-IT1801-52.